OpenStudy (anonymous):

Using complete sentences, explain what might happen if the order of operations was used to solve a literal equation.

OpenStudy (anonymous):

Define literal.

OpenStudy (anonymous):

A literal equation is an equation with multiple variables. Solving a literal equation means to solve for one of these specific variables.

OpenStudy (anonymous):

Like 7x+6z+4y = 14...? x=?

OpenStudy (anonymous):

example: E=mc^2, solve for c. E=mc2 -- ---- m m E/m=c^2 c^2=e/m √c^2=√e/m c=√e/m

OpenStudy (anonymous):

You wouldn't be able to solve the equation, because you would still have the variable... many times, you have to use addition or subtraction to get the term you want to solve for on one side. ?

OpenStudy (anonymous):

im lost... sorry

OpenStudy (saifoo.khan):